# Extracting energy from plunging region of a Kerr-Taub-NUT black hole by magnetic reconnection

### v1

@@ Abstract
We have studied the energy extraction from a Kerr-Taub-NUT black hole via magnetic reconnections occurring in the plunging region. Our results show that the gravitomagnetic charge suppresses the energy extraction process through magnetic reconnection and reduces the corresponding extraction efficiency, which is opposite to those of the black hole spin and the magnetization parameter. Finally, we treat the energy extraction process through magnetic reconnection as a mechanism to revisit the problem on the observed jet power and radiative efficiency of GRS 1915+105. Our results show that the allowed black hole parameter region originating from the jet power has an intersection with the region from the radiative efficiency. This means that with this mechanism related to magnetic reconnection the Kerr-Taub-NUT metric can simultaneously explain the observed jet power and radiative efficiency for GRS 1915+105, which is not explained by other mechanisms in previous studies.
@@ 1 Introduction
Extracting energy from rotating black holes has been an intriguing topic since it was proposed by Penrose. In the Penrose process, a particle splits into two parts in the ergosphere. One of parts possessing negative energy viewed from infinity falls into the black hole. The resultant part escapes from the ergosphere to infinity with more energy than the original one beacuse of the energy conservation law. This implies that one can extract the rotational energy of a Kerr black hole through this process, until the black hole’s mass is reduced to its irreducible value [1]. However, initiation mechanisms are lacked in the original Penrose process [2]. One plausible mechanism is the superradiance of a massive bosonic field because the superradiant cloud can be formed with a characteristic spectrum surrounding the black hole [3]. The Blandford-Znajek process [4] is another mechanism to extract the energy of a rotating black hole, where the magnetic field lines originating from accretion flows near the black hole are assumed to intersect the event horizon as well as they possess the angular velocity smaller than that of the horizon. The Blandford-Znajek (BZ) mechanism has been extensively investigated since it is expected as a promsing engine to generate relativistic jets [5, 6, 7].
@@
Another energy extraction mechanism has been investigated in the past twenty decades, which is based on magnetic reconnections in plasmoides [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. In this mechanism, the magnetic reconnection plays as the trigger to split a plasmoide into two parts as in Penrose process. Koide and Arai discussed firstly the possibility of extracting energy from black holes by this mechanism [12]. Recently, the plasmoid instability within the reconnection layer and the generation of expelled plasmoids are found to yield fast reconnection and fast Alfvén outflows [13, 14, 15, 16]. Subsequently, the fast reconnection model has been applied to study the energy extraction through the plasmoids around black holes [17, 18], which show that the energy extraction rate could be comparable to that of the BZ process. Due to its astrophysical significance, the energy extraction by magnetic reconnections have been studied in different types of spacetimes [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. However, most of above studies on energy extraction via magnetic reconnections have been mainly focused on circular accretion flows outside the ISCO. Due to instability of circular orbits within the ergosphere, the accretion flows inside the ISCO plunge toward the event horizon with spiral-shape orbits. Therefore, energy extraction from a Kerr black hole has been studied through magnetic reconnection within the plunging region [32]. Since plunging region is near the black hole event horizon, the behavior of the ejected plasmoids produced from magnetic reconnections and the corresponding energy extraction carry significantly features related to the background spacetime, which could provide a potenial way to test theories of gravity. Therefore, it is necessary to further study energy extraction by magnetic reconnections from the plunging region in other background spacetimes of rotating black holes.
@@
Kerr-Taub-NUT (KTN) metric is an important stationary and axisymmetric metric, which satisfies Einstein field equations with gravitomagnetic monopole and dipole moments [33, 34, 35, 36, 37, 38, 39]. It owns three parameters: the mass $M$, the spinning parameter $a$ and the NUT charge $l$. The charge $l$ plays the role of a magnetic mass inducing a topology in the Euclidean section. Moreover, due to the existence of the NUT charge, KTN spacetime is asymptotically non-flat and there exist conical singularities on the axis of symmetry [40]. According to Misner [41], this singularity can be avoided by introducing a periodic time coordinate. However, the periodicity condition inevitably gives rise to closed timelike curves (CTCs) in this spacetime, thereby leading to violations of causality and the incompleteness of geodesics [42]. Fortunately, these problems can be resolved by giving up the time periodicity condition as well as interpreting the conical singularities as “a linear source of pure angular momentum” [40, 43, 44, 45]. Although the KTN spacetime exhibits certain unfavorable physical characteristics, it remains highly attractive for exploring various physical phenomena in general relativity. The black hole shadow is found to increase with the NUT charge [46]. Moreover, the NUT charge affects particle accelerations in the spacetime and further modifies the restrict conditions of the rotation parameter $a$ when the arbitrarily high center-of-mass energy appears in the collision of two particles [47]. The effects of the NUT charge on the inner-most stable circular orbits and the perfect fluid disk around the black hole have also been studied in Refs.[48, 49], respectively. The parameter constraint on the Kerr-Taub-NUT black hole is carried out with the observed jet power and radiative efficiency [50]. The observational possibilities of NUT charge are also analyzed in terms of the shifting of spectral lines from quasars, supernovae and active galactic nuclei [51]. These results may help to probe the gravitomagnetic masses in astronomical observation in the future. The main purpose of this paper is to study energy extraction via magnetic reconnection in the plunging region near a Kerr-Taub-NUT black hole and probe corresponding effects from NUT charge.
@@
The paper is organized as follows. In Sec. 2, we briefly introduce the Kerr-Taub-NUT spacetime and magnetic reconnection. In Sec. 3, we analyze energy extraction by magnetic reconnections in the Kerr-Taub- NUT spacetime and probe how NUT charge affect the corresponding energy extraction. In Sec. 4, we treat the energy extraction process through magnetic reconnection as a mechanism to revisit the problem on the observed jet power and radiative efficiency of GRS 1915+105. Finally, we end with a summary.
@@ 2 The magnetic reconnection in the Kerr-Taub-NUT spacetime
The Kerr-Taub-NUT metric describes the gravitational field of a black hole with the mass parameter $M$, the gravitomagnetic charge $l$, and the rotational parameter $a$. Its form reads [52]
$ds^{2}=$displaystyle-- $\frac{1}{\Sigma}(\Delta-a^{2}\sin^{2}\theta)dt^{2}+\Sigma\mathopen{}\mathclose{{}\left(\frac{1}{\Delta}dr^{2}+d\theta^{2}}\right)+\frac{1}{\Sigma}\mathopen{}\mathclose{{}\left[(\Sigma+a\chi)^{2}\sin^{2}\theta-\chi^{2}\Delta}\right]d\phi^{2}$ (2.1)
$+$ $\frac{2}{\Sigma}\mathopen{}\mathclose{{}\left(\Delta\chi-a(\Sigma+a\chi)\sin^{2}\theta}\right)d\phi dt\ ,$
@@
By setting $\Delta=0$, we obtain the radius of outer horizon for the black hole
@@
The radius of the outer ergoregion surface is
@@
At two poles ($\theta=0$ or $\pi$), the outer ergoregion surface touches the event horizon. While on the equatorial plane ($\theta=\pi/2$), the ergoregion radius $r_{e}|_{\theta=\frac{\pi}{2}}=M+\sqrt{M^{2}+l^{2}}$ depends only on the NUT parameter and the mass, and has no connection to the black hole spin. Setting the denominator of the metric component $g_{tt}$ to zero, we obtain the location of singularity
$r=0\,\,\,\text{and}\,\,\,\theta=\cos^{-1}(-l/a)\,.$
When $l>a$, the spacetime becomes free of singularity. The Kerr-Taub-NUT spacetime describes a black hole when $|a|\leq\sqrt{M^{2}+l^{2}}$ and a naked singularity when $a>\sqrt{M^{2}+l^{2}}$.
Let us now probe magnetic reconnections in the Kerr-Taub-NUT spacetime. As in [32], we adopt the perfect fluid approximation for the ejected plasmoids and the bulk plasma in the plunging region near the black hole. In this approximation, the stress-energy tensor of the plasma is
@@
For the plasma moving on the equatorial plane, the flow is along timelike geodesics without $\theta$-dependence. Therefore, we have $U_{\mu}dx^{\mu}=-Edt+U_{r}(r)dr+Ld\phi$, where $E$ and $L$ is the conserved energy and angular momentum along the streamline. To properly describe the occurrences within the fluid during magnetic reconnection, one must project the fluid 4-velocity into the normal tetrad of the zero-angular momentum observers
@@
$\hat{e}^{\,\mu}_{(t)}=\frac{1}{\alpha}(\partial_{t}^{\,\mu}+\omega^{\phi}\partial_{\phi}^{\,\mu})\,,\quad\hat{e}^{\,\mu}_{(r)}=\frac{1}{h_{r}}\partial_{r}^{\,\mu}\,,\quad\hat{e}^{\,\mu}_{(\theta)}=\frac{1}{h_{\theta}}\partial_{\theta}^{\,\mu}\,,\quad\hat{e}^{\,\mu}_{(\phi)}=\frac{1}{h_{\phi}}\partial_{\phi}^{\,\mu}\,.$ (2.10)
@@
As in [32], here we focus only on the fast reconnection model within the ideal MHD framework [18]. In this model, the thermal pressure is assumed to be negligible relative to the magnetic pressure, which generally exists in highly magnetized accretion systems, including magnetically arrested disks [53, 54]. Moreover, the magnetic field lines are assumed to have opposite directions upon crossing the equatorial plane and magnetic reconnection occurs within a much smaller region [55] near the dominant reconnection $p$-point [56] so that reconnection can be regarded to take place at the $p$-point on the global scale. Due to spontaneous perturbations, the two sets of anti-parallel magnetic field lines $\vec{B}_{0}$ curved toward each other, which yields that the local magnetic pressure reduces so further push the anti-parallel magnetic lines towards each other more until they touch at the current sheet. Then reconnection occurs and its energy makes the plasmoid burst into two parts, ejecting along the direction of the upstream magnetic field line $\vec{B}_{0}$. In a word, the reconnection plays a similar role of ”bomb” in the Penrose process. In the fast reconnection model, the size of current sheet can be comparable to the mean free paths of ions [8]. The current sheet can be simply characterized by the geometry index as
@@
where $d$ and $L$ denote the half-width and half-length of the current sheet, respectively. Assuming that the plasmoids do not slow down after leaving the current sheet and using the MHD equilibrium, the magnitude of their ejection velocities can be estimated as [32]
@@
which has a peak-value at ${\sf g}\simeq 0.49$. As in [32], in the cases of high local-magnetization, magnetic reconnection can be assumed to occur at the maximum local-reconnection rate with the geometric index ${\sf g}=0.49$.
The magnetic field of a black hole is generally thought to be antiparallel adjacent to the equatorial plane, which is crucial for the occurrence of magnetic reconnection. For simplification, we consider a magnetic field which at infinity approaches a poloidal configuration, which means its radial component is non-zero. For plasma moving along non-circular plunging streamlines, we can assume the fluid is idea-MHD because the conductivity of the plasma is nearly zero outside the current sheet so that the magnetic field lines are frozen in flowing plasma[57, 58]. From the equation $F^{\mu\nu}U_{\mu}=0$, one can find that for a stationary and axisymmetric fluid configuration composed of zero-conductivity plasma in the Kerr-Taub-NUT spacetime the global magnetic field can be expressed as
$B_{0}^{t}=-\frac{\Psi}{\sqrt{-g}}\bigg{(}U_{r}+\frac{U^{\phi}}{U^{r}}L\bigg{)}\,,\quad B_{0}^{r}=-\frac{\Psi}{\sqrt{-g}}E\,,\quad B_{0}^{\phi}=-\frac{\Psi}{\sqrt{-g}}\frac{U^{\phi}}{U^{r}}E\,,$ (2.15)
@@
$B_{0}^{[1]}=\frac{\Psi}{\sqrt{-g}\,\hat{U}}\sqrt{g_{rr}g_{\phi\phi}}\,\omega^{\phi}\,,\quad B_{0}^{[3]}=\frac{\Psi}{\sqrt{-g}\,\hat{U}U^{r}}(\hat{\gamma}_{s}E-\alpha)\,,$ (2.16)
where $\hat{U}=\sqrt{\mathopen{}\mathclose{{}\left(\hat{U}^{(r)}}\right)^{2}+\mathopen{}\mathclose{{}\left(\hat{U}^{(\phi)}}\right)^{2}}$. In the fluid’s rest frame, the plasmoids are ejected along the magnetic field line. Therefore, the ejection direction of plasmoids can be described by the angle between the magnetic field line and $e_{[3]}$,
@@
where $\gamma_{out}$ is the Lorentz factor measured in the fluid’s rest frame and the sign “$\pm$” denotes two plasmoids towards opposite directions. In the reconnection process, one can neglect the magnetic energy because in the current sheet most of them is converted to the plasma kinetic energy. Furthermore, with the incompressible sphere approximation, one can find that the proper enthalpy density in the plasmoid is equal to that of the bulk plasma at the $p$-point [12, 32] and the enthalpy is $\Delta H=\omega\Delta V$, where $\Delta V$ is a small proper volume of the plasmoid. The energy per enthalpy at infinity for ejected plasmoids is
@@
ϵ±=subscriptitalic-$\epsilon_{\pm}=$ $\,\,\alpha\hat{\gamma}_{s}\gamma_{out}\bigg{[}\big{(}1+\beta^{\phi}\hat{v}^{(\phi)}_{s}\big{)}\pm v_{out}\bigg{(}\hat{v}_{s}+\beta^{\phi}\frac{\hat{v}^{(\phi)}_{s}}{\hat{v}_{s}}\bigg{)}\cos{\xi}\mp v_{out}\beta^{\phi}\frac{\hat{v}^{(r)}_{s}}{\hat{\gamma}_{s}\hat{v}_{s}}\sin{\xi}\bigg{]}$
@@
The spacetime directly affects the energy per enthalpy of plasmoid through the factors $\alpha$ and $\beta^{\phi}$, but also indirectly through the fluid velocity and the magnetic field angle.

### v2

@@ Abstract
We have studied the energy extraction from a Kerr-Taub-NUT black hole via magnetic reconnection occurring in the plunging region. Our results show that the gravitomagnetic charge suppresses the energy extraction process through magnetic reconnection and reduces the corresponding extraction efficiency, which is opposite to the effects of the black hole spin and the magnetization parameter. Finally, we treat the energy extraction process through magnetic reconnection as a mechanism to revisit the problem of the observed jet power and radiative efficiency of GRS 1915+105. Our results show that the allowed black hole parameter region originating from the jet power has an intersection with the region from the radiative efficiency. This means that with this mechanism related to magnetic reconnection the Kerr-Taub-NUT metric can simultaneously explain the observed jet power and radiative efficiency for GRS 1915+105, which is not explained by other mechanisms in previous studies.
@@ 1 Introduction
Extracting energy from rotating black holes has been an intriguing topic since it was proposed by Penrose. In the Penrose process, a particle splits into two parts in the ergosphere. One of the parts, possessing negative energy viewed from infinity, falls into the black hole. The resultant part escapes from the ergosphere to infinity with more energy than the original one because of the law of energy conservation. This implies that one can extract the rotational energy of a Kerr black hole through this process, until the black hole’s mass is reduced to its irreducible value [1]. However, initiation mechanisms are lacking in the original Penrose process [2]. One plausible mechanism is the superradiance of a massive bosonic field because a superradiant cloud can be formed with a characteristic spectrum surrounding the black hole [3]. The Blandford-Znajek process [4] is another mechanism for extracting the energy of a rotating black hole, where the magnetic field lines originating from accretion flows near the black hole are assumed to intersect the event horizon and possess an angular velocity smaller than that of the horizon. The Blandford-Znajek (BZ) mechanism has been extensively investigated since it is expected as a promising engine for generating relativistic jets [5, 6, 7].
@@
Another energy extraction mechanism has been investigated over the past twenty years, which is based on magnetic reconnections in plasmoids [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 12, 21, 22, 23]. Magnetic reconnection refers to the sudden rearrangement of magnetic fields induced by the local interaction of magnetic field lines with opposing polarities. Such topological changes of magnetic fields yield a rapid release of magnetic energy into thermal and kinetic energy of plasma outflows. One of the famous classical magnetic reconnection models is the Sweet-Parker model [9, 10]. In this model, magnetic reconnection proceeds with the formation of an elongated diffusion region. However, its predicted reconnection rate is excessively low, failing to account for the observed phenomena in most astrophysical systems. To address this challenge, Petschek [11] proposed a fast magnetic reconnection model in which the diffusion region is localized and the outflow is bounded by a pair of standing slow mode shocks. Within this reconnection geometry, the thickness and length of the diffusion region reside on the microscopic scale, resulting in a large aspect ratio for the diffusion region. This, in turn, produces a faster reconnection rate that is sufficient to account for the observed time scales of solar flares and geomagnetic substorms. Moreover, the Petschek model means that slow mode shocks could play an important role in reconnection. However, numerical simulations have shown that Petschek reconnection cannot form in two-dimensional resistive-MHD simulations if the resistivity is uniform [24, 25]. To further understand fast magnetic reconnection, a series of theoretical investigations and numerical simulations have been performed both in collisionless [26] and collisional plasmas [27, 28], and different mechanisms are analyzed from the perspectives of plasma instabilities [29, 30, 31, 32], anomalous resistivity[33, 34, 35], turbulence [36, 37, 38, 39, 40, 41] and Hall effects [42, 43, 44, 45, 46]. Recently, based on magnetic reconnection, Koide and Arai first discussed the possibility of extracting energy from black holes [47]. The fast reconnection model has also been applied to study the energy extraction through the plasmoids around black holes [48, 49], which shows that the energy extraction rate could be comparable to that of the BZ process. Due to its astrophysical significance, the energy extraction by magnetic reconnections has been studied in different types of spacetimes [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. However, most of the above studies on energy extraction via magnetic reconnection have been mainly focused on circular accretion flows outside the ISCO. This is because the circular orbits within the ISCO are unstable and the accretion flows inside the ISCO plunge toward the event horizon with spiral-shape orbits. Recently, energy extraction from a Kerr black hole has been studied through magnetic reconnection within the plunging region [63]. Since the plunging region is near the black hole event horizon, the behavior of the ejected plasmoids produced from magnetic reconnection and the corresponding energy extraction carry significant features related to the background spacetime, which could provide a potential way to test theories of gravity. Therefore, it is necessary to further study energy extraction by magnetic reconnections from the plunging region in other background spacetimes of rotating black holes.
@@
The Kerr-Taub-NUT (KTN) metric is an important stationary and axisymmetric metric, which satisfies Einstein field equations with gravitomagnetic monopole and dipole moments [64, 65, 66, 67, 68, 69, 70]. It owns three parameters: the mass $M$, the spin parameter $a$ and the NUT charge $l$. The charge $l$ plays the role of a magnetic mass inducing a topology in the Euclidean section. Moreover, due to the existence of the NUT charge, KTN spacetime is asymptotically non-flat and there exist conical singularities on the axis of symmetry [71]. According to Misner [72], this singularity can be avoided by introducing a periodic time coordinate. However, the periodicity condition inevitably gives rise to closed timelike curves (CTCs) in this spacetime, thereby leading to violations of causality and the incompleteness of geodesics [73]. Fortunately, these problems can be resolved by giving up the time periodicity condition as well as interpreting the conical singularities as “a linear source of pure angular momentum” [71, 74, 75, 76]. Although the KTN spacetime exhibits certain unfavorable physical characteristics, it remains highly attractive for exploring various physical phenomena in general relativity. The black hole shadow is found to increase with the NUT charge [77]. Moreover, the NUT charge affects particle accelerations in the spacetime and further modifies the constraints on the rotation parameter $a$ when the arbitrarily high center-of-mass energy appears in the collision of two particles [78]. The effects of the NUT charge on the inner-most stable circular orbits and the perfect fluid disk around the black hole have also been studied in Refs.[79, 80], respectively. The parameter constraint on the Kerr-Taub-NUT black hole is carried out with the observed jet power and radiative efficiency [81]. The observational possibilities of NUT charge are also analyzed in terms of the shifting of spectral lines from quasars, supernovae and active galactic nuclei [82]. These results may help to probe the gravitomagnetic masses in astronomical observation in the future. The main purpose of this paper is to study energy extraction via magnetic reconnection in the plunging region near a Kerr-Taub-NUT black hole and to probe corresponding effects from NUT charge.
@@
The paper is organized as follows. In Sec. 2, we briefly introduce the Kerr-Taub-NUT spacetime and magnetic reconnection. In Sec. 3, we analyze energy extraction by magnetic reconnection in the Kerr-Taub-NUT spacetime and probe how the NUT charge affects the corresponding energy extraction. In Sec. 4, we treat the energy extraction process through magnetic reconnection as a mechanism to revisit the problem on the observed jet power and radiative efficiency of GRS 1915+105. Finally, we conclude with a summary.
@@ 2 The magnetic reconnection in the Kerr-Taub-NUT spacetime
The Kerr-Taub-NUT metric describes the gravitational field of a black hole with the mass parameter $M$, the gravitomagnetic charge $l$, and the rotational parameter $a$. Its form reads [64, 65, 66, 67, 68, 69, 70]
$ds^{2}$ $=$ $-\frac{1}{\Sigma}(\Delta-a^{2}\sin^{2}\theta)dt^{2}+\frac{\Sigma}{\Delta}dr^{2}+\Sigma d\theta^{2}+\frac{1}{\Sigma}\mathopen{}\mathclose{{\left[(\Sigma+a\chi)^{2}\sin^{2}\theta-\chi^{2}\Delta}}\right]d\phi^{2}$ (2.1)
$+$ $\frac{2}{\Sigma}\mathopen{}\mathclose{{\left[\Delta\chi-a(\Sigma+a\chi)\sin^{2}\theta}}\right]d\phi dt\ ,$
@@
By setting $\Delta=0$, one can obtain the radius of the outer horizon for the black hole
@@
Correspondingly, the radius of the outer ergoregion surface is given by
@@
At the two poles ($\theta=0$ or $\pi$), the outer ergoregion surface touches the event horizon. On the equatorial plane ($\theta=\pi/2$), however, the ergoregion radius $r_{e}|_{\theta=\frac{\pi}{2}}=M+\sqrt{M^{2}+l^{2}}$ depends only on the NUT parameter and the mass, and has no connection to the black hole spin. By setting the denominator of the metric component $g_{tt}$ to zero, we obtain the location of the singularity
$r=0,\quad\quad\text{and}\quad\quad\theta=\cos^{-1}(-l/a)\,.$
When $l>a$, the spacetime becomes free of singularities. The Kerr-Taub-NUT spacetime describes a black hole when $|a|\leq\sqrt{M^{2}+l^{2}}$ and a naked singularity when $a>\sqrt{M^{2}+l^{2}}$.
Let us now probe magnetic reconnection in the Kerr-Taub-NUT spacetime. As in [63], we adopt the perfect fluid approximation for the ejected plasmoids and the bulk plasma in the plunging region near the black hole. In this approximation, the stress-energy tensor of the plasma is
@@
For the plasma moving on the equatorial plane, the flow is along timelike geodesics without $\theta$-dependence. Therefore, one has $U_{\mu}dx^{\mu}=-Edt+U_{r}(r)dr+Ld\phi$, where $E$ and $L$ are the conserved energy and angular momentum along the streamline. To properly describe the physical processes within the fluid during magnetic reconnection, one must project the fluid 4-velocity into the normal tetrad of the zero-angular momentum observers (ZAMOs)
@@
$\hat{e}^{\,\mu}_{(t)}=\frac{1}{\alpha}(\partial_{t}^{\,\mu}+\omega^{\phi}\partial_{\phi}^{\,\mu})\,,\quad\quad\hat{e}^{\,\mu}_{(r)}=\frac{1}{h_{r}}\partial_{r}^{\,\mu}\,,\quad\quad\hat{e}^{\,\mu}_{(\theta)}=\frac{1}{h_{\theta}}\partial_{\theta}^{\,\mu}\,,\quad\quad\hat{e}^{\,\mu}_{(\phi)}=\frac{1}{h_{\phi}}\partial_{\phi}^{\,\mu}\,.$ (2.10)
@@
As in [63], here we focus only on the fast reconnection model within the ideal MHD framework [49]. In this model, the thermal pressure is assumed to be negligible relative to the magnetic pressure, which generally exists in highly magnetized accretion systems, including magnetically arrested disks [83, 84]. Moreover, the magnetic field lines are assumed to have opposite directions upon crossing the equatorial plane and magnetic reconnection occurs within a much smaller region [85] near the dominant reconnection $p$-point [86] so that reconnection can be regarded to take place at the $p$-point on the global scale. Due to spontaneous perturbations, the two sets of anti-parallel magnetic field lines $\vec{B}_{0}$ curve toward each other, which which causes the local magnetic pressure to reduce and further push the anti-parallel magnetic lines towards each other more until they touch at the current sheet. Then reconnection occurs and its energy makes the plasmoid burst into two parts, ejecting along the direction of the upstream magnetic field line $\vec{B}_{0}$. In a word, the reconnection plays a role similar to that of “bomb” in the Penrose process. In the fast reconnection model, the size of the current sheet can be comparable to the mean free paths of ions [8]. The current sheet can be simply characterized by the geometry index as
@@
where $d$ and $L$ denote the half-width and half-length of the current sheet, respectively. Assuming that the plasmoids do not slow down after leaving the current sheet and using the MHD equilibrium condition, the magnitude of their ejection velocities can be estimated as [63]
@@
which has a peak-value at ${\sf g}\simeq 0.49$. As in [63], in the cases of high local-magnetization, magnetic reconnection can be assumed to occur at the maximum local reconnection rate with the geometric index ${\sf g}=0.49$.
The magnetic field of a black hole is generally thought to be antiparallel adjacent to the equatorial plane, which is crucial for the occurrence of magnetic reconnection. For the plasma moving along non-circular plunging streamlines, we can assume the fluid is stationary, axisymmetric, and ideal-MHD because the resistivity of the plasma is nearly zero outside the current sheet so that the magnetic field lines are frozen in flowing plasma [87, 88]. For such a magnetic field, there exists a field line angular velocity $\Omega_{B}$ characterizing the rotation of magnetic field lines. For simplification, we here adopt the assumption of a zero field line angular velocity $\Omega_{B}=0$ which significantly reduces the complexity of calculations. From the equation $F^{\mu\nu}U_{\mu}=0$, one can find that for a stationary and axisymmetric fluid configuration composed of a zero-resistivity plasma in the Kerr-Taub-NUT spacetime, the global magnetic field can be expressed as
$B_{0}^{t}=-\frac{\Psi}{\sqrt{-g}}\bigg(U_{r}+\frac{U^{\phi}}{U^{r}}L\bigg),\quad\quad\quad B_{0}^{r}=-\frac{\Psi}{\sqrt{-g}}E,\quad\quad\quad B_{0}^{\phi}=-\frac{\Psi}{\sqrt{-g}}\frac{U^{\phi}}{U^{r}}E,$ (2.15)
@@
$B_{0}^{[1]}=\frac{\Psi}{\sqrt{-g}\,\hat{U}}\sqrt{g_{rr}g_{\phi\phi}}\,\omega^{\phi}\,,\quad\quad\quad B_{0}^{[3]}=\frac{\Psi}{\sqrt{-g}\,\hat{U}U^{r}}(\hat{\gamma}_{s}E-\alpha)\,$, (2.16)
where $\hat{U}=\sqrt{\mathopen{}\mathclose{{\left(\hat{U}^{(r)}}}\right)^{2}+\mathopen{}\mathclose{{\left(\hat{U}^{(\phi)}}}\right)^{2}}$. In the fluid’s rest frame, the plasmoids are ejected along the magnetic field lines. Therefore, the ejection direction of plasmoids can be described by the angle between the magnetic field lines and $e_{[3]}$,
@@
where $\gamma_{out}$ is the Lorentz factor measured in the fluid’s rest frame and the sign $\pm$” denotes two plasmoids towards opposite directions. In the reconnection process, one can neglect the magnetic energy because in the current sheet most of it is converted to the plasma’s kinetic energy. Furthermore, with the incompressible sphere approximation, one can find that the proper enthalpy density in the plasmoids is equal to that of the bulk plasma at the $p$-point [47, 63] and the enthalpy is $\Delta H=\omega\Delta V$, where $\Delta V$ is a small proper volume of the plasmoids. The energy per enthalpy at infinity for ejected plasmoids is
@@
$\epsilon_{\pm}=$ $\,\,\alpha\hat{\gamma}_{s}\gamma_{out}\bigg[\big(1+\beta^{\phi}\hat{v}^{(\phi)}_{s}\big)\pm v_{out}\bigg(\hat{v}_{s}+\beta^{\phi}\frac{\hat{v}^{(\phi)}_{s}}{\hat{v}_{s}}\bigg)\cos{\xi}\mp v_{out}\beta^{\phi}\frac{\hat{v}^{(r)}_{s}}{\hat{\gamma}_{s}\hat{v}_{s}}\sin{\xi}\bigg]$
@@
The spacetime affects the energy per unit enthalpy of the plasmoid not only directly through the factors $\alpha$ and $\beta^{\phi}$, but also indirectly through the fluid velocity and the magnetic field angle.